OpenStudy (anonymous):

Algebra 2 help, Simplify the expression; sqrt(-1) / (2 -4i)-(5+3i)

OpenStudy (mathstudent55):

Is this the problem? \(\dfrac{\sqrt{-1}}{(2 - 4i) - (5 + 3i)} \)

OpenStudy (anonymous):

I can attach a image or draw it out if needed

OpenStudy (mathstudent55):

That means it's not what I asked?

OpenStudy (mathstudent55):

Can you show what the problem is supposed to be if I didn't understand it correctly?

OpenStudy (anonymous):

\[\frac{ \sqrt{-1} }{ (2-4i)-(5+3i) }\]

OpenStudy (mathstudent55):

Ok. Start by simplifying \( \sqrt{-1} \)

OpenStudy (anonymous):

i

OpenStudy (mathstudent55):

Good.

OpenStudy (mathstudent55):

Now you need to do the subtraction of complex numbers in the denominator. Drop the first set of parentheses because it is unnecessary. Then distribute the negative to the left of the second set of parentheses.

OpenStudy (anonymous):

2-4i - (-)5-3i

OpenStudy (anonymous):

-3-7i

OpenStudy (mathstudent55):

\(\dfrac{ \sqrt{-1} }{ (2-4i)-(5+3i) }\) \(= \dfrac{ i}{ 2-4i-(5+3i) }\) Be careful with the distribution of the negative: \(= \dfrac{ i}{ 2-4i-5-3i }\)

OpenStudy (mathstudent55):

The 5 inside the parentheses is positive, so it becomes negative, or -5. The 3i is positive, so it becomes -3i.

OpenStudy (anonymous):

So would it be -7-7i? or -3-7i?

OpenStudy (mathstudent55):

Now you combine like terms in the denominator.

OpenStudy (mathstudent55):

\(= \dfrac{ i}{ -3-7i }\)

OpenStudy (anonymous):

So now you use the conjugate to eliminate the i, right

OpenStudy (mathstudent55):

Now we need one more step, because we're not supposed to leave i in the denominator.

OpenStudy (mathstudent55):

Correct, but it's to eliminate i from the denominator. There can be i in the numerator.

OpenStudy (anonymous):

\[(\frac{ i }{ -3-7i }) \times (\frac{ 3+7i }{ 3+7i })\]

OpenStudy (mathstudent55):

Be careful with the complex conjugate. The only sign that changes is the sign on the term with i. \(= \dfrac{ i}{ -3 - 7i } \times \dfrac{-3 + 7i}{-3 + 7i}\)

OpenStudy (anonymous):

Oh wow, I think that's where i've been going wrong

OpenStudy (mathstudent55):

The conjugate of a + bi is a - bi, not -a - bi. In fact what you did, by changing both signs, you are using the opposite (also called the additive inverse) of the complex number. That is not the complex conjugate.

OpenStudy (mathstudent55):

Now you can multiply out the numerators and denominators.

OpenStudy (anonymous):

\[\frac{ -3i-7 }{ 58 }\]

OpenStudy (mathstudent55):

\(= \dfrac{ -3i + 7i^2 }{(-3)^2 - (7i)^2}\) \(= \dfrac{ -3i - 7 }{9 - 49(-1)}\) \(= \dfrac{-7 - 3i}{9 + 49} \) \(=\dfrac{-7 - 3i}{58} \) \(=-\dfrac{7}{58} - \dfrac{3}{58}i\)

OpenStudy (mathstudent55):

You are correct.

OpenStudy (anonymous):

Thank you so much!